OpenStudy (anonymous):

Linda had 84 fliers to post around town. Last week, she posted 1/3 of them. This week, she posted 1/4 of the remaining films. How many fliers has she still not posted?

OpenStudy (anonymous):

Fliers posted last week: 1/3 x 84 = 28 Remaining fliers: 84 - 28 = 56 Fliers posted this week: 1/4 x 56 = 14 Fliers not posted: 56 - 14 = 42